OpenStudy (anonymous):

Find the sum of:

OpenStudy (anonymous):

\[2\sqrt{6}\] and \[\sqrt{24}\]

OpenStudy (anonymous):

simplify sqrt(24) first

OpenStudy (anonymous):

A. \[4\sqrt{6}\] B. \[2\sqrt{30}\] C. \[2\sqrt{6}+4\sqrt{3}\]

OpenStudy (anonymous):